# A Method for On‑Orbit Calibration of the VLAST‑P Electromagnetic Calorimeter

### v2

@@ Abstract
Very Large Area Gamma-ray Space Telescope Pathfinder (VLAST-P), serving as the technology validation satellite for VLAST mission, is designed to observe high-energy solar bursts on orbit. The CsI electromagnetic calorimeter (ECAL) is one of the key sub-detectors of VLAST-P. In order to study the on-orbit energy calibration method of ECAL, the simulation of VLAST-P was carried out based on Geant4. The results indicated an energy resolution better than 10 % in the 0.1–5 GeV range, while a dedicated minimum-ionization-particle (MIP) calibration method was developed both to ensure accurate energy reconstruction and to monitor detector stability throughout the in-orbit calibration period.
@@ 1 Space-Based High-Energy Gamma-Ray Observatories
High-energy gamma-ray astronomy has evolved through two main approaches: space-based satellites that directly detect gamma rays and ground-based instruments that observe the particle showers these rays produce in the atmosphere. Ground-based techniques are generally more effective for very high-energy (VHE) gamma rays, as they are less constrained by size and cost, whereas space-based satellites are better suited for observations in the X-ray and high-energy (HE, 30 MeV-30 GeV) ranges [Gamma-ray_Astronomy].
@@
The Compton Gamma Ray Observatory (CGRO), one of NASA’s gamma ray observatories in the 1990s, was equipped with four gamma-ray telescopes that collectively covered a broad energy range from the keV to GeV scale. For gamma rays with energies above 10 MeV, where pair production becomes the dominant interaction mechanism, CGRO’s high-energy observations were primarily performed by the Energetic Gamma Ray Experiment Telescope (EGRET) [Thompson2008]. EGRET was designed as a high-energy gamma-ray detector featuring a multilayer spark chamber system for tracking, triggered by a scintillator coincidence system, which was sensitive in the energy range 20 MeV to 30 GeV. Energy measurement was carried out using a total absorption shower counter based on NaI(Tl) crystal, with an overall thickness of 8 radiation length. EGRET achieved an energy resolution of 20–25% across most of its sensitive energy range [EGRETCalib].
@@
The Astrorivelatore Gamma ad Immagini LEggero (AGILE) was developed by the Italian Space Agency (ASI) for high-energy astrophysics research. The AGILE payload included four active detectors that together covered a wide energy range, from hard X-rays to gamma rays. Among them, the gamma-ray imaging detector was made up of a segmented anti-coincidence system, a silicon tracker, and a non-imaging CsI(Tl) mini-calorimeter [Vercellone2024pgh].
@@
The Large Area Telescope (LAT) onboard the Fermi Gamma-ray Space Telescope is a pair-conversion gamma-ray detector with a wide field of view (2.4 sr) and an energy range from below 20 MeV to more than 300 GeV [FermiLAT]. It consists of a silicon-strip tracker, a hodoscopic CsI(Tl) calorimeter, and a segmented anti-coincidence detector. The LAT is designed to measure photon direction, energy, and arrival time while rejecting cosmic-ray background. Since its launch, it has enabled all-sky surveys, source localization, spectral measurements, and studies of gamma-ray bursts, diffuse backgrounds, and potential dark matter signals.
@@
In the near future, The Very Large Area gamma-ray Space Telescope (VLAST) [2022AcASn..63...27F, Pan_2024] is a proposed mission designed to observe gamma rays across a broad energy spectrum, from MeV to TeV. It will utilize two primary detection methods: Compton scattering and electron-positron pair production. The telescope will operate in low Earth orbit, conducting a detailed survey of gamma-ray sources. Its instrumentation includes: an anti-coincidence detector (ACD) to filter out background events, a tracker detector that also functions as a low-energy calorimeter and a high energy imaging calorimeter (HEIC) for precise measurements.
@@ 2 Very Large Area Gamma-ray Space Telescope Pathfinder
Figure 1 shows the schematics of the whole detector system. It maintains the design of VLAST detector system. Its detector system comprises three main sub-detector systems: ACD, tracker and electro-magnetic calorimeter (ECAL). The ACD of VLAST-P is designed with two primary technical objectives: to achieve effective separation between charged particles and gamma rays (with a charged particle detection efficiency exceeding 99.9%), maintain optimal performance at high event rates (up to $10\,\mathrm{kHz/m^{2}}$). Tracker detector is built from two main components: a 15 mm thick CsI converter and a silicon tracker. The CsI converter is placed upon silicon tracker after ACD. It acts as both a photon converter that converts the incoming photons to a pair of electron and positron and an active calorimeter with radiation length of around $0.8\,X_{0}$. The benefit of utilizing a sensitive detector as converter instead of a traditional tungsten converter is that one could collect the energy deposited in CsI which completes the energy measurements. It could be of vital importance in low energy particle reconstructions. Following the converter, the silicon tracker consists of six silicon micro-strip layers arranged in an alternating x-y configuration. Each pair of orthogonal strips provides 2D coordinate measurements. This design characterizes the electron-positron pairs from photon conversion with a spatial resolution better than $50\,\mathrm{um}$ [AZZARELLO2016378].
The main sub-detector for particle energy measurement is the electromagnetic calorimeter (ECAL). The detector is optimized to cover an energy range of $100\,\mathrm{MeV}$ to $5\,\mathrm{GeV}$, although the ECAL design is limited by the area and weight available on the satellite. It is built with 25 CsI(Tl) crystal bars arranged in a $5\times 5$ array, as shown in the lower part of figure 1. Thallium-doped cesium iodide is chosen as the sensitive material because of its high light yield and good uniformity.
The size of each CsI(Tl) crystal is $60\,\mathrm{mm}\times 60\,\mathrm{mm}\times 200\,\mathrm{mm}$. With a radiation length ($X_{0}$) of $1.86\,\mathrm{cm}$ for CsI(Tl), the longitudinal depth of the ECAL corresponds to approximately $10.8\,X_{0}$, which is sufficient to contain a large portion of the electromagnetic showers from GeV photons. The crystals are wrapped with $\mathrm{TiO_{2}}$ reflective films to fully collect the scintillation photons. The crystals are separated by a mechanical gap of $1\,\mathrm{mm}$; thus, the total active volume is approximately $300\,\mathrm{mm}\times 300\,\mathrm{mm}\times 200\,\mathrm{mm}$.
@@
The required readout energy range is designed to cover the energy range up to $5\,\mathrm{GeV}$. To measure incident photons up to $5\,\mathrm{GeV}$, simulations indicate a maximum energy deposition of approximately $4.1\,\mathrm{GeV}$ in a single crystal, which determines the upper limit of the measurement range. At the low-energy end, the energy threshold of single crystal in simulation is set to $3\,\mathrm{MeV}$; therefore, the electronic noise must be controlled to meet the threshold. As a result, the readout system is required to cover a wide dynamic range from the $3\,\mathrm{MeV}$ up to the $4.1\,\mathrm{GeV}$ saturation level.
@@ 3.1 Geant4 Simulation Framework
In this work, the Geant4 toolkit [ALLISON2016186, AGOSTINELLI2003250] is employed to perform detailed Monte Carlo simulations of particle interactions within the VLAST-P payload. Geant4 is a comprehensive framework for modeling the transport and interaction of high-energy particles with detector materials, incorporating both electromagnetic and hadronic processes across a broad energy range. Its modular architecture allows for the accurate implementation of geometry, material composition, and detector response, ensuring a realistic description of the experimental setup. For the present study, the simulation focuses on the response of the CsI(Tl)-based electromagnetic calorimeter to cosmic-ray protons, helium nuclei, and solar $\gamma$-rays, which constitute the dominant sources of background and calibration signals in orbit. The use of Geant4 thus enables quantitative evaluation of detector performance, including energy deposition, shower development, and calibration stability, providing critical input for both event reconstruction and on-orbit calibration strategies.
@@ 3.2 Detector Simulation Setup
Figure 2 shows the exploded view of our simulated detector system drawn with front-end tool JSROOT [JSROOT]. The components are displayed in separate layers to illustrate the internal structure. As shown in the picture from left to right, the two outermost plates (shown in light green) are the ACD layers, which are positioned at the front to effectively distinguish charged particles from photons. Next to the ACD, four CsI(Tl) plates (shown in yellow) are placed before the six silicon tracker layers (depicted in blue and red layers). Specifically, four CsI(Tl) plates serve as the converter while the six silicon microstrip plates (used to measure x and y coordinates) provide position information with 3 oriented along the x-axis and 3 along the y-axis. Each of the two adjacent plates constitutes a layer that provides position information for the hit. Finally, on the far right, the ECAL is visualized as the $5\times 5$ array of purple crystal bars, which are used to measure the energy of particles entering the crystals. Detailed mechanical and electronic components are also visible: the small red squares attached to the crystal ends are APDs, while the purple frame-like structures correspond to the silicone rubber pads designed for shock absorption. Inclusion of these fine details ensures a more realistic representation of the detector geometry.
@@ 3.3 Performance and Validation
By employing a full Geant4-based simulation, the intrinsic energy response of the calorimeter to monoenergetic photons was evaluated. Gamma rays with energies ranging from 50 MeV to 5 GeV were simulated at normal incidence to the detector surface. The primary particles were emitted from a planar source of size $150\penalty 10000\ \mathrm{mm}\times 150\penalty 10000\ \mathrm{mm}$, positioned 10 mm away from the center of the ACD’s front surface. The particles were incident perpendicularly to the detector surface. A total of $1\times 10^{5}$ events were generated to ensure sufficient statistical precision for each incident energy point. The energy of each event is reconstructed by summing the energies of adjacent crystals surrounding the one with the maximum deposition. The VLAST-P tracker consists of two components: an active converter layer, where incident gamma rays are expected to undergo pair production, and a silicon tracking section that records the subsequent electron-positron trajectories with high spatial resolution. The tracker serves two essential functions: (i) photon conversion, by providing an active medium in which gamma rays can convert into an electron–positron pair, and (ii) direction reconstruction, by accurately measuring the pair trajectories to infer the incoming photon’s direction. In the context of the ECAL simulation, tracker information is used to identify whether the photon has converted in the tracker; events that convert upstream are either rejected or treated separately, ensuring that the energy response to calorimeter is evaluated only for photons entering the ECAL as intended. As illustrated in figure 3(a), the energy distribution at each incident energy point is fitted with a Crystal Ball function [PhysRevD.34.711] to extract the mean reconstructed energy and the full width at half maximum (FWHM). The energy resolution $R$ is calculated from the FWHM of the energy peak as follows:
@@
Figure 3: Performance of the VLAST-P ECAL: estimated energy linearity and energy resolution.
The performance of the VLAST-P ECAL was evaluated in terms of both energy linearity and resolution for incident gamma rays. As shown in figure 3(b), the measured energy deposited in the CsI(Tl) scintillator exhibits a nearly proportional response to the true incident energy over the range from $20\,\mathrm{MeV}$ to $5\,\mathrm{GeV}$, demonstrating good energy linearity with less than 10% deviations. Correspondingly, the energy resolution, illustrated in figure 3(c), remains around 5% across a wide energy range for gamma rays normally incident on the detector. This is due to the high stopping power of the CsI(Tl) crystals, which absorb most of the incident energy. Consequently, only a small fraction of energy leakage occurs even at higher energies, ensuring accurate energy reconstruction.
The resolution curve shows a slight dip around 500 MeV, which corresponds to an optimal balance between sufficient energy deposition and minimal fluctuation in shower containment. At this point, the electromagnetic shower is largely contained within the detector volume, and stochastic fluctuations are relatively suppressed. As the incident energy increases beyond this point, the resolution gradually worsens due to increasing longitudinal and transverse shower leakage, which leads to larger fluctuations in the reconstructed energy and thus a degradation in resolution.
@@ 4.1 Simulation Methodology
Backtracing is generally easier and more efficient than forward tracing, as it launches from the observation point and avoids sampling the vast incident phase space [osti_4784046]. A comprehensive global geomagnetic backtracing database, named GeoMagFilter, has been constructed in Ref. [CHEN20255450]. It was generated using specialized software that simulates particle trajectories via an eighth-order Runge-Kutta numerical integration method within the International Geomagnetic Reference Field (IGRF-13) model. The database is constructed with three parameters: particle rigidity, zenith angle, and azimuth angle. Particle rigidity is defined as momentum divided by the charge quantity of the particle to represent how difficult it is for a particle to be deflected by a magnetic field. The latter two parameters can be calculated based on geographic longitude and latitude, which are typically accessible as basic orbital information of the satellite. The sampling grid spans longitudes from $0^{\circ}$ to $360^{\circ}$ in $10^{\circ}$ intervals and latitudes from $75^{\circ}$ to $75^{\circ}$ in $5^{\circ}$ steps, covering all possible orbital positions of both the International Space Station (ISS) and the China Space Station (CSS). Directional angles are sampled uniformly on a spiral in $4\pi$ steradians with an angular spacing of 0.03 radians, resulting in approximately 13,963 distinct directions. A binary value is assigned to indicate whether a particle can successfully trace back beyond the geomagnetic field (1 for successful backtracing, 0 for failure, e.g., when the trajectory terminates within $10^{5}$ km).

### v3

@@ Abstract
Very Large Area Gamma-ray Space Telescope Pathfinder (VLAST-P), serving as the technology validation satellite for VLAST mission, is designed to observe high-energy solar bursts on orbit. The CsI electromagnetic calorimeter (ECAL) is one of the key sub-detectors of VLAST-P. In order to study the on-orbit energy calibration method of ECAL, the simulation of VLAST-P was carried out based on Geant4. The results indicated an energy resolution better than 10 % in the 0.1–5 GeV range, while a dedicated minimum-ionization-particle (MIP) calibration method was developed both to ensure accurate energy reconstruction and to monitor detector stability throughout the on-orbit calibration period.
@@ 1 Space-Based High-Energy Gamma-Ray Observatories
High-energy gamma-ray astronomy has evolved through two main approaches: space-based satellites that directly detect gamma rays and ground-based instruments that observe the particle showers these rays produce in the atmosphere. Ground-based techniques are generally more effective for very high-energy (VHE) gamma rays, as they are less constrained by size and cost, whereas space-based satellites are better suited for observations in the X-ray and high-energy (HE, 30 MeV-30 GeV) ranges [10].
@@
The Compton Gamma Ray Observatory (CGRO), one of NASA’s gamma ray observatories in the 1990s, was equipped with four gamma-ray telescopes that collectively covered a broad energy range from the keV to GeV scale. For gamma rays with energies above 10 MeV, where pair production becomes the dominant interaction mechanism, CGRO’s high-energy observations were primarily performed by the Energetic Gamma Ray Experiment Telescope (EGRET) [15]. EGRET was designed as a high-energy gamma-ray detector featuring a multilayer spark chamber system for tracking, triggered by a scintillator coincidence system, which was sensitive in the energy range 20 MeV to 30 GeV. Energy measurement was carried out using a total absorption shower counter based on NaI(Tl) crystal, with an overall thickness of 8 radiation length. EGRET achieved an energy resolution of 20–25% across most of its sensitive energy range [16].
@@
The Astrorivelatore Gamma ad Immagini LEggero (AGILE) was developed by the Italian Space Agency (ASI) for high-energy astrophysics research. The AGILE payload included four active detectors that together covered a wide energy range, from hard X-rays to gamma rays. Among them, the gamma-ray imaging detector was made up of a segmented anti-coincidence system, a silicon tracker, and a non-imaging CsI(Tl) mini-calorimeter [17].
@@
The Large Area Telescope (LAT) onboard the Fermi Gamma-ray Space Telescope is a pair-conversion gamma-ray detector with a wide field of view (2.4 sr) and an energy range from below 20 MeV to more than 300 GeV [5]. It consists of a silicon-strip tracker, a hodoscopic CsI(Tl) calorimeter, and a segmented anti-coincidence detector. The LAT is designed to measure photon direction, energy, and arrival time while rejecting cosmic-ray background. Since its launch, it has enabled all-sky surveys, source localization, spectral measurements, and studies of gamma-ray bursts, diffuse backgrounds, and potential dark matter signals.
@@
In the near future, the Very Large Area gamma-ray Space Telescope (VLAST) [8, 14] is a proposed mission designed to observe gamma rays across a broad energy spectrum, from MeV to TeV. It will utilize two primary detection methods: Compton scattering and electron-positron pair production. The telescope will operate in low Earth orbit, conducting a detailed survey of gamma-ray sources. Its instrumentation includes: an anti-coincidence detector (ACD) to filter out background events, a tracker detector that also functions as a low-energy calorimeter and a high energy imaging calorimeter (HEIC) for precise measurements.
@@ 2 Very Large Area Gamma-ray Space Telescope Pathfinder
Figure 1 shows the schematics of the whole detector system. It maintains the design of VLAST detector system. Its detector system comprises three main sub-detector systems: ACD, tracker and electro-magnetic calorimeter (ECAL). The ACD consists of two layers of plastic scintillators: the upper layer is a single plate ($400\,\mathrm{mm}\times 400\,\mathrm{mm}$), while the lower layer is segmented into four smaller tiles ($200\,\mathrm{mm}\times 200\,\mathrm{mm}$) arranged in an overlapping configuration to minimize dead zones. The signals are read out by PMTs coupled with wavelength-shifting fibers. This design is driven by two primary technical objectives: to achieve effective separation between charged particles and gamma rays (with a charged particle detection efficiency exceeding 99.9%), maintain optimal performance at high event rates (up to $10\,\mathrm{kHz/m^{2}}$). The VLAST-P tracker consists of two main components: a $15\,\mathrm{mm}$ thick CsI converter ($\sim 0.8\,X_{0}$) and a silicon tracking section. It serves a dual purpose: acting as an active medium for photon conversion into electron-positron pairs, and reconstructing the incident direction by measuring the pair trajectories with high precision. Unlike traditional passive tungsten converters, the active CsI layer measures the energy deposited during conversion. This capability allows for complete energy reconstruction, which is critically important for low-energy particles.Following the converter, the silicon tracker consists of six silicon micro-strip layers arranged in an alternating x-y configuration. Each pair of orthogonal strips provides 2D coordinate measurements. By adopting the silicon strip detector technology successfully verified in the DAMPE mission [6], this design is able to characterize the electron-positron pairs from photon conversion with a spatial resolution better than $50\,\mu\mathrm{m}$.
The main sub-detector for particle energy measurement is the electromagnetic calorimeter (ECAL). Given the stringent constraints on satellite weight and dimensions, the ECAL is optimized to balance performance with available resources. It is composed of 25 CsI(Tl) crystal bars arranged in a $5\times 5$ array, as shown in the lower part of figure 1. Thallium-doped cesium iodide (CsI(Tl)) is selected as the sensitive material due to its high light yield and excellent uniformity. Each crystal measures $60\,\mathrm{mm}\times 60\,\mathrm{mm}\times 200\,\mathrm{mm}$. With a radiation length ($X_{0}$) of $1.86\,\mathrm{cm}$ for CsI(Tl), the longitudinal depth of approximately $10.8\,X_{0}$ is sufficient to contain a significant fraction (about 93% for 1 GeV photons) of the electromagnetic shower. Transversely, with a Molière radius of $3.57\,\mathrm{cm}$, the $300\,\mathrm{mm}\times 300\,\mathrm{mm}$ cross-section ensures effective containment of the shower core. The crystals are wrapped with $\mathrm{TiO_{2}}$ reflective films to fully collect the scintillation photons. The crystals are separated by a mechanical gap of $1\,\mathrm{mm}$; thus, the total active volume is approximately $300\,\mathrm{mm}\times 300\,\mathrm{mm}\times 200\,\mathrm{mm}$.
@@
The calorimeter is designed to measure incident photons with energies up to 5 GeV. For a 5 GeV photon, simulations indicate that the total energy deposited in the full ECAL approaches 5 GeV. The maximum energy deposition observed in a single crystal is approximately 4.1 GeV, which determines the required dynamic range of the front-end readout electronics, since each channel must accommodate the largest possible energy deposited in an individual crystal. At the low-energy end, the energy threshold of single crystal in simulation is set to $3\,\mathrm{MeV}$; therefore, the electronic noise must be controlled to meet the threshold. As a result, the readout system is required to cover a wide dynamic range from $3\,\mathrm{MeV}$ up to approximately 4.1 GeV
@@ 3.1 Geant4 Simulation Framework
In this work, the Geant4 toolkit [3, 1] is employed to perform detailed Monte Carlo simulations of particle interactions within the VLAST-P payload. Geant4 is a comprehensive framework for modeling the transport and interaction of high-energy particles with detector materials, incorporating both electromagnetic and hadronic processes across a broad energy range. Its modular architecture allows for the accurate implementation of geometry, material composition, and detector response, ensuring a realistic description of the experimental setup. For the present study, the simulation focuses on the response of the CsI(Tl)-based electromagnetic calorimeter to cosmic-ray protons, helium nuclei, and solar $\gamma$-rays, which constitute the dominant sources of background and calibration signals in orbit. The use of Geant4 thus enables quantitative evaluation of detector performance, including energy deposition, shower development, and calibration stability, providing critical input for both event reconstruction and on-orbit calibration strategies.
@@ 3.2 Detector Simulation Setup
Figure 2 shows the exploded view of our simulated detector system drawn with front-end tool JSROOT [12]. The components are displayed in separate layers to illustrate the internal structure. As shown in the picture from left to right, the two outermost plates (shown in light green) are the ACD layers, which are positioned at the front to effectively distinguish charged particles from photons. Next to the ACD, four CsI(Tl) plates (shown in yellow) are placed before the six silicon tracker layers (depicted in blue and red layers). Specifically, four CsI(Tl) plates serve as the converter while the six silicon microstrip plates (used to measure x and y coordinates) provide position information with 3 oriented along the x-axis and 3 along the y-axis. Each of the two adjacent plates constitutes a layer that provides position information for the hit. Finally, on the far right, the ECAL is visualized as the $5\times 5$ array of purple crystal bars, which are used to measure the energy of particles entering the crystals. Detailed mechanical and electronic components are also visible: the small red squares attached to the crystal ends are APDs, while the purple frame-like structures correspond to the silicone rubber pads designed for shock absorption. Inclusion of these fine details ensures a more realistic representation of the detector geometry.
@@ 3.3 Performance and Validation
By employing a full Geant4-based simulation, the intrinsic energy response of the calorimeter to monoenergetic photons was evaluated. Gamma rays with energies ranging from 50 MeV to 5 GeV were simulated at normal incidence to the detector surface. The primary particles were emitted from a planar source of size $150\penalty 10000\ \mathrm{mm}\times 150\penalty 10000\ \mathrm{mm}$, positioned 10 mm away from the center of the ACD’s front surface. The particles were incident perpendicularly to the detector surface. A total of $1\times 10^{5}$ events were generated to ensure sufficient statistical precision for each incident energy point. The energy of each event is reconstructed by summing the energy deposits in the crystal with the maximum signal and its neighboring crystals, together with the energy deposited in the active converter. As shown in figure 3, this combined energy reconstruction is important because a non-negligible fraction of the incident energy, particularly for low-energy gamma rays, is deposited in the converter before the electromagnetic shower develops in the ECAL. If only ECAL signals are used, the reconstructed energy is systematically underestimated also leading to a visible low-energy tail at 50 and 100 MeV. By including the energy deposited in the converter, this bias is largely corrected, and the reconstructed energy distribution shifts closer to unity, improving the accuracy of the low-energy measurement. In the context of the ECAL simulation, tracker information is used to identify whether the photon has converted in the tracker; to evaluate the energy reconstruction performance for the pair-production channel, we specifically select events where the photon converts into an electron-positron pair within the CsI converter. The total reconstructed energy is defined as the sum of the energy deposited in the active CsI converter and the energies measured by ECAL crystals ($E_{\mathrm{rec}}=E_{\mathrm{CsI}}+E_{\mathrm{ECAL}}$). This ensures that the contribution of the active converter to the total energy measurement is properly included in the evaluation. As illustrated in figure 4(a), the energy distribution at each incident energy point is fitted with a Crystal Ball function [9] to extract the mean reconstructed energy and the full width at half maximum (FWHM). The energy resolution $R$ is calculated from the FWHM of the energy peak as follows:
@@
Figure 3: Comparison of reconstructed energy using only the ECAL (left) versus including the active converter’s deposit (right).
@@
Figure 4: Performance of the VLAST-P ECAL: estimated energy linearity and energy resolution. Only events with photon conversion process are selected for analysis.
The performance of the VLAST-P ECAL was evaluated in terms of both energy linearity and resolution for incident gamma rays. As shown in figure 4(b), the measured energy deposited in the CsI(Tl) scintillator exhibits a nearly proportional response to the true incident energy over the range from $50\,\mathrm{MeV}$ to $5\,\mathrm{GeV}$, demonstrating good energy linearity with less than 5% deviations. The resolution curve, illustrated in figure 4(c), shows a sharp decrease in the low-energy range, dropping from approximately $10\%$ to about $5\%$ at $1000\,\mathrm{MeV}$. Above $1\,\mathrm{GeV}$, the improvement slows down, and the resolution stabilizes at a plateau of approximately $6\%$ up to $5\,\mathrm{GeV}$, indicating stable performance across the high-energy range. Consequently, only a small fraction of energy leakage occurs even at higher energies, ensuring accurate energy reconstruction.
The longitudinal containment of the ECAL is defined by its thickness of 10.8 $X_{0}$. Monte Carlo simulations with 100 MeV and 1 GeV gamma rays indicate that approximately 93% of the incident energy is contained within this 20-cm-long CsI crystal. The remaining leakage is attributed to the finite detector depth, and future reconstruction improvements will include a dedicated leakage correction for high-energy events. The transverse containment is governed by the Molière radius of CsI, which is 3.57 cm, and it is generally accepted that 95% of the shower energy is contained within two Molière radii. When the shower maximum is located within the central $3\times 3$ crystal array, the ECAL geometry is sufficient to contain nearly all transverse shower energy deposition.
@@ 4.1 Simulation Methodology
Backtracing is generally easier and more efficient than forward tracing, as it launches from the observation point and avoids sampling the vast incident phase space [13]. A comprehensive global geomagnetic backtracing database, named GeoMagFilter, has been constructed in Ref. [7]. It was generated using specialized software that simulates particle trajectories via an eighth-order Runge-Kutta numerical integration method within the International Geomagnetic Reference Field (IGRF-13) model. The database is constructed with three parameters: particle rigidity, zenith angle, and azimuth angle. Particle rigidity is defined as momentum divided by the charge quantity of the particle to represent how difficult it is for a particle to be deflected by a magnetic field. The latter two parameters can be calculated based on geographic longitude and latitude, which are typically accessible as basic orbital information of the satellite. The sampling grid spans longitudes from $0^{\circ}$ to $360^{\circ}$ in $10^{\circ}$ intervals and latitudes from $75^{\circ}$ to $75^{\circ}$ in $5^{\circ}$ steps, covering all possible orbital positions of both the International Space Station (ISS) and the China Space Station (CSS). Directional angles are sampled uniformly on a spiral in $4\pi$ steradians with an angular spacing of 0.03 radians, resulting in approximately 13,963 distinct directions. A binary value is assigned to indicate whether a particle can successfully trace back beyond the geomagnetic field (1 for successful backtracing, 0 for failure, e.g., when the trajectory terminates within $10^{5}$ km).